William drove 5 hours at the average speed of 54 miles an hour. For the first 2 hours, he drove 45 miles an hour. What was his average for the last 3 hours?
Answer by mananth(16946)

William drove 5 hours at the average speed of 54 miles an hour. For the first 2 hours, he drove 45 miles an hour. What was his average for the last 3 hours?
time driven = 5 hours
54 miles / hour
Distance traveled = 54 * 5= 270 miles
2 hours @ 45 mph
distance traveled = 90 miles
270-90= 180 miles is the second leg distance
180/3 = 60 mph
Next 3 hours he drove at 60 mph